# Text-Speech Language Models with Improved Cross-Modal Transfer by Aligning Abstraction Levels

### v1

@@ Appendix B Datasets details
Text FineWeb-Edu (Ben Allal et al., 2024) 4B 0.7
Cosmopedia-v2 (Ben Allal et al., 2024) 4B 0.15
Python-Edu (Ben Allal et al., 2024) 2B 0.08
FineMath (Liu et al., 2024) 2B 0.06
Speech LibriSpeech (Panayotov et al., 2015) 67M (960 hours)
LibriLight (Kahn et al., 2020) 3.7B (53k hours)
SWC (Baumann et al., 2019) 32M (1k hours)
Tedlium (Hernandez et al., 2018) 0.1B (1.6k hours)
People (Galvez et al., 2021) 0.5B (7k hours)
Vox Populi (Wang et al., 2021) 1.6B (24k hours)
sTinyStories (Cuervo & Marxer, 2024) 4.8B (72k hours)
Interleaved text-speech LibriHeavy (Kang et al., 2024) 313M 3.1B (50k hours) 0.37
sTinyStories (Cuervo & Marxer, 2024) 800M 4.8B (72k hours) 0.53
SWC (Baumann et al., 2019) 3.6M 26M (800 hours) 0.1
@@ C.1 Intrinsic dimensionality and subspace overlap
The intrinsic dimensionality and subspace overlap are estimated using five batches of 10k representations each, totaling 50k samples. We compute each metric per batch and report average and deviations. To obtain each representation, we extract random subsequences of 20 words and use the final representation in the sequence as the sequence representation. For the intrinsic dimensionality text samples are randomly drawn from FineWeb-Edu, while speech samples are taken from the full set of speech datasets. For the subspace overlap we use paired samples from the sTinyStories dataset.
@@
To estimate the intrinsic dimensionality we use the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) (Denti et al., 2022) implementation in dadapy (Glielmo et al., 2022) and follow the procedure described by Cheng et al. (2025).
@@ C.2 Word segmentation
We apply a peak detector to the sequences of last-layer dynamic weights, $\boldsymbol{\omega}^{(L)}$, using SciPy’s find_peaks tool. Performance is evaluated as a binary prediction task, where a peak indicates the prediction of a boundary at a given position. We allow a tolerance of one token to account for noise in the boundary annotations. As in standard binary prediction tasks, we use recall, precision, and the F1-score as performance metrics. Additionally, we use the R-value (Räsänen et al., 2009), which penalizes trivial over-segmentation solutions. To optimize performance, we tune the prominence parameter of the peak detector over a grid $(0,0.15]$ with steps of $0.01$ so as to maximize the R-value, following Kreuk et al. (2020).
@@ Appendix D Additional results
Extended benchmark results, including the smaller baselines and other speech LMs from the literature, are presented in Table 7. Figure 6 illustrates the scaling of several metrics for the baseline and SmolTolk models.
Figures 7, 8, and 9 provide additional results on the analysis of multi-layer pooling for SmolTolk-2B, SmolTolk-400M, and SmolTolk-150M, respectively.
@@
GSLM (Lakhotia et al., 2021) 100M Ø 54.2 Ø 66.6 Ø Ø Ø 53.3 Ø Ø Ø
AudioLM (Borsos et al., 2023) 150M Ø 64.7 Ø Ø Ø Ø Ø Ø Ø
Hassid et al. (2023) cold-init 1.3B 1.3B 10.8B Ø 56.5 Ø Ø Ø Ø Ø Ø Ø
TWIST (Hassid et al., 2023) 1.3B 1.3B 10.8B Ø 57.0 Ø 70.6 Ø Ø Ø 52.4 Ø Ø Ø
TWIST (Hassid et al., 2023) 7B 7B 36B Ø 59.0 Ø 74.1 Ø Ø Ø 55.3 Ø Ø Ø
TWIST (Hassid et al., 2023) 13B 13B 36B Ø 59.2 Ø 76.4 Ø Ø Ø 55.4 Ø Ø Ø
Cuervo & Marxer (2024) best 823M 82B Ø 61.3 Ø 78.0 Ø Ø Ø 56.7 Ø Ø Ø
SyllableLM (Baade et al., 2025) 300M 1.2B Ø 63.7 Ø 75.4 Ø Ø Ø Ø Ø Ø
AlignSLM (Lin et al., 2024) 7B 7B Ø 62.3 Ø 86.8 Ø Ø Ø 61.1 Ø Ø Ø
Slam (scaled) (Maimon et al., 2025) 358M 16.7B Ø 61.1 Ø 84.2 Ø Ø Ø 61.3 Ø Ø Ø
@@
LAST (Turetzky & Adi, 2024) $\sim$390M 56.8
Moshi (Défossez et al., 2024) 7.7B 2.1T 58.8 83.0 60.8 49.8 / 54.3
Zeng et al. (2024) 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
Zeng et al. (2024) 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3

### v2

@@ Appendix B Datasets details
Text FineWeb-Edu [Allal et al., 2025] 4B 0.7
Cosmopedia-v2 [Allal et al., 2025] 4B 0.15
Python-Edu [Allal et al., 2025] 2B 0.08
FineMath [Liu et al., 2024] 2B 0.06
Speech LibriSpeech [Panayotov et al., 2015] 67M (960 hours)
LibriLight [Kahn et al., 2020] 3.7B (53k hours)
SWC [Baumann et al., 2019] 32M (1k hours)
Tedlium [Hernandez et al., 2018] 0.1B (1.6k hours)
People [Galvez et al., 2021] 0.5B (7k hours)
Vox Populi [Wang et al., 2021] 1.6B (24k hours)
sTinyStories [Cuervo and Marxer, 2024] 4.8B (72k hours)
Interleaved text-speech LibriHeavy [Kang et al., 2024] 313M 3.1B (50k hours) 0.37
sTinyStories [Cuervo and Marxer, 2024] 800M 4.8B (72k hours) 0.53
SWC [Baumann et al., 2019] 3.6M 26M (800 hours) 0.1
@@ C.1 Intrinsic dimensionality and subspace overlap
The intrinsic dimensionality and subspace overlap are estimated using five batches of 10k representations each, totaling 50k samples. We compute each metric per batch and report average and standard deviations across batches. To obtain each representation, we extract random subsequences of 20 words and use the final representation in the sequence as the sequence representation. For the intrinsic dimensionality text samples are randomly drawn from FineWeb-Edu, while speech samples are taken from the full set of speech datasets. For the subspace overlap we use paired samples from the sTinyStories dataset.
@@
Intrinsic Dimensionality. To estimate the intrinsic dimensionality we use the Generalized Ratios Intrinsic Dimension Estimator (GRIDE) [Denti et al., 2022] implementation in dadapy [Glielmo et al., 2022] and follow the procedure described by Cheng et al. [2025].
Cross-modal subspace overlap.
Let $\mathbf{X}\!\in\!\mathbb{R}^{n\times d_{X}}$ be the target-modality embeddings (e.g. text) and $\mathbf{Y}\!\in\!\mathbb{R}^{n\times d_{Y}}$ the source-modality embeddings (e.g. speech), both mean-centred. We measure how much of $\mathbf{Y}$ can be expressed inside the $k$-dimensional principal subspace of $\mathbf{X}$:
  1. (i)
    PCA on $\mathbf{X}$. Compute the top-$k$ eigenvectors $\mathbf{U}^{(k)}_{X}\in\mathbb{R}^{d_{X}\times k}$ of the covariance of $\mathbf{X}$ and the associated variance fraction $v_{X}(k)=\sum_{i=1}^{k}\lambda^{(X)}_{i}$.
  2. (ii)
    Project & reconstruct $\mathbf{Y}$.
    $\widehat{\mathbf{Y}}\;=\;\mathbf{Y}\,\mathbf{U}^{(k)}_{X}\,\mathbf{U}^{(k)\!\top}_{X}$.
  3. (iii)
    Fraction of variance in $\mathbf{Y}$ explained.
[…]
A score of $1$ indicates that the $\mathbf{Y}$ variance lies entirely within the dominant $\mathbf{X}$ span; lower values signal weaker linear alignment between the two modalities. In all experiments we use $k=50$.
@@ C.2 Word segmentation
We apply a peak detector to the sequences of last-layer dynamic weights, $\boldsymbol{\omega}^{(L)}$, using SciPy’s find_peaks tool. Performance is evaluated as a binary prediction task, where a peak indicates the prediction of a boundary at a given position with a tolerance of one token to account for noise in the boundary annotations. As in standard binary prediction tasks, recall, precision, and the F1-score are used as performance metrics. Additionally, we use the R-value [Räsänen et al., 2009], which penalizes trivial over-segmentation solutions. To optimize performance, we tune the prominence parameter of the peak detector over a grid $(0,0.15]$ with steps of $0.01$ so as to maximize the R-value, following Kreuk et al. [2020].
@@ Appendix D Additional results
D.1 Cross-modal subspace overlap for larger models
We compute the cross-modal representation space alignment score described in Section 5.2 for our largest models—namely, the 1.7B baseline, SmolTolk-2B, and Spirit LM, a much larger model (7B parameters trained on over 170B tokens). Unlike Figure 3, where we compared models with the same representation dimensionality, here the models have different dimensions, and thus may exhibit different spectral tails. To account for this, we normalize the overlap score by the fraction of variance captured by the $\mathbf{Y}$ subspace:
$\frac{f(k)}{v_{Y}(k)}.$
In Figure 5, a normalized score of $1$ indicates that the dominant variance of $\mathbf{Y}$ lies entirely within the dominant subspace of $\mathbf{X}$.
Several observations are worth noting: larger models with early fusion/fission—such as the baseline and Spirit LM—exhibit increased alignment in the middle layers. We hypothesize that, by this point, these models have developed sufficiently high-level representations for some text-learned functions at those layers to become useful. This likely contributes to the improved cross-modal performance observed in the 1.7B baseline (Figure 2). As suggested in Valeriani et al. [2023], Cheng et al. [2025], models are expected to reach a peak of compositionality, after which some layers may be dedicated to processing high-level features. We further expect that deeper models have more capacity for such processing (rather than immediate refinement toward next-token prediction), which may help explain Spirit LM’s strong performance. Greater depth gives speech representations more opportunities to undergo composition and approximate the abstraction level needed to exploit text-learned functions in later layers.
Nonetheless, alignment remains substantially lower than in SmolTolk-2B. As observed at smaller scales, early fusion models struggle to leverage text-learned functionality due to insufficient composition in early layers. Similarly, the absence of late multi-level fission limits abstraction and may force these models to use their final layers for extracting speech-predictive features, thereby preventing or disrupting the use of next-word predictive knowledge encoded in those layers. Given these observations, we expect that at a similar scale to Spirit LM, a model with our architecture would outperform it.
Figure 5: Cross-modal alignment score in our largest models and Spirit LM
D.2 Low-level speech information across layers
Figure 7 shows the cross-entropy loss for affine probes trained to predict the speech token $w_{i}$ from the representation $\mathbf{c}_{i}^{(l)}$ at each layer $l$ of the models in Table 2. This measures how much low-level speech information is linearly preserved across the model’s layers. Notably, models without multi-level fission tend to preserve more of such information throughout layers.
Figure 6: Forgetting of the current speech token across layers.
D.3 SmolTolk cross-modal generations
As illustrated in Table 7, SmolTolk can generate semantically consistent content when prompted with either speech or text tokens. Audio prompts are taken from unseen LibriSpeech test-clean samples, while text prompts were manually crafted to resemble those in Table 1 of Nguyen et al. [2024], for comparison purposes.
Setup Prompt Generation
$\toS T$he world is [swt] [Hu24][Hu12][Hu341]...[Hu500]
n too small for all the ideas of mine…
[…]
n This cat was very special because it could…
[…]
n Illustration: the lemon [swt]
is a fruit that grows on trees it is round and yellow and tastes sour…
[…]
n The Earth has undoubtedly entered upon a new orbit, but she is now encountering a [swt]
new phase of its history to be followed by a new and marvelous system of life upon The Earth…
Table 7: SmolTolk-2B cross-modal generations. The [swt] token denotes a modality switch in the generation process. HuBERT speech tokens are denoted as [Hu]. SmolTolk-2B displays coherent and consistent semantic cross-modal abilities.
D.4 SmolTolk + ASR and TTS
Model Params. Tokens # Shots LS clean LS other
$downarrow ASR\downarrow TTS\downarrow$
Spirit LM [Nguyen et al., 2024] 7B $\sim$175B 10 6.0 6.7 11.0 7.9
SmolTolk-2B+ASR+TTS 2B 16B 0 8.0 12.0 19.2 14.5
SmolTolk-2B+ASR+TTS 2B 16B 10 7.8 10.0 17.4 12.9
Table 8: Automatic Speech Recognition (ASR) and Text-to-Speech (TTS) performance on the LibriSpeech (LS) dataset. ASR scores are reported as Word Error Rate (WER, %), and TTS scores as Character Error Rate (CER, %). SmolTolk-2B is evaluated with a maximum context length of 2048 tokens. Few-shot examples are sampled from the corresponding LibriSpeech development sets.
Model Params. Tokens tStoryCloze sStoryCloze
[…]
SmolTolk-2B 2B 16B 92.6 87.5 83.9 86.0 73.2 60.0 64.0 63.4
SmolTolk-2B+ASR+TTS 2B 16B 92.3 87.0 83.1 86.0 72.6 60.6 62.2 62.6
Table 9: StoryCloze evaluation results comparing SmolTolk-2B and the more multitask variant trained with a mixture of interleaved, pure speech, pure text, ASR, and TTS samples.
A common application of TSLMs is traditional speech processing tasks such as ASR and TTS. While our focus is on text LLM-like linguistic capabilities rather than these applications, we perform an additional experiment where, in addition to the tasks described in Section 3.1—speech modeling, text modeling, and interleaved text-speech modeling—we also include ASR and TTS training. This follows similar multi-task setups in prior work [Chou et al., 2023, Nguyen et al., 2024, Zeng et al., 2024, Défossez et al., 2024], and allows us to assess how expanding the task mix affects model performance.
In this setup, each training batch allocated one-twelfth of samples to ASR, one-twelfth to TTS, one-sixth to interleaved samples, and the remaining two-thirds to speech and text modeling. Table 8 reports zero-shot and few-shot results on LibriSpeech. For TTS evaluation, we follow Nguyen et al. [2024], using Whisper Medium [Radford et al., 2022] to transcribe speech synthesized by the HiFiGAN vocoder from Hassid et al. [2023]. Despite being 3.5× smaller and trained on 11× less data than Spirit LM, SmolTolk-2B+ASR+TTS achieves competitive ASR and TTS performance. Results improve with more in-context examples, highlighting the model’s ability to adapt in context.
Finally, to assess whether adding ASR and TTS compromises cross-modal reasoning, we evaluate on the StoryCloze benchmark. As shown in Table 9, performance across other tasks is only minimally affected.
D.5 Scaling of other metrics
@@
Figure 7 illustrates the scaling of several metrics for the baseline and SmolTolk models.
D.6 Extended benchmark table
Extended benchmark results, including the smaller baselines and other speech LMs from the literature, are presented in Table 10.
@@ D.6 Extended benchmark table
GSLM [Lakhotia et al., 2021] 100M Ø 54.2 Ø 66.6 Ø Ø Ø 53.3 Ø Ø Ø
AudioLM [Borsos et al., 2023] 150M Ø 64.7 Ø Ø Ø Ø Ø Ø Ø
Hassid et al. [2023] cold-init 1.3B 1.3B 10.8B Ø 56.5 Ø Ø Ø Ø Ø Ø Ø
TWIST [Hassid et al., 2023] 1.3B 1.3B 10.8B Ø 57.0 Ø 70.6 Ø Ø Ø 52.4 Ø Ø Ø
TWIST [Hassid et al., 2023] 7B 7B 36B Ø 59.0 Ø 74.1 Ø Ø Ø 55.3 Ø Ø Ø
TWIST [Hassid et al., 2023] 13B 13B 36B Ø 59.2 Ø 76.4 Ø Ø Ø 55.4 Ø Ø Ø
Cuervo and Marxer [2024] best 823M 82B Ø 61.3 Ø 78.0 Ø Ø Ø 56.7 Ø Ø Ø
SyllableLM [Baade et al., 2025] 300M 1.2B Ø 63.7 Ø 75.4 Ø Ø Ø Ø Ø Ø
AlignSLM [Lin et al., 2024] 7B 7B Ø 62.3 Ø 86.8 Ø Ø Ø 61.1 Ø Ø Ø
Slam (scaled) [Maimon et al., 2025] 358M 16.7B Ø 61.1 Ø 84.2 Ø Ø Ø 61.3 Ø Ø Ø
@@
Spirit LM [Nguyen et al., 2024] 7B $\sim$175B 73.3 59.7 95.8 90.5 78.6 94.3 74.0 66.3 64.7 71.7 37.7 / 39.0
LAST [Turetzky and Adi, 2024] $\sim$390M 56.8
Moshi [Défossez et al., 2024] 7.7B 2.1T 58.8 83.0 60.8 49.8 / 54.3
GLM-4-Voice [Zeng et al., 2024] 1.5B 1.5B 1T 77.5 81.4 90.1 55.4 58.6 64.0
GLM-4-Voice [Zeng et al., 2024] 9B 9B 1T 83.0 85.0 93.6 62.4 63.2 76.3
@@
D.7 More multi-layer pooling samples and weights details
Figures 8, 9, and 10 provide additional results on the analysis of multi-layer pooling for SmolTolk-2B, SmolTolk-400M, and SmolTolk-150M, respectively.